OpenStudy (raphaelll):

Simplify the following expressions. Give your answers in power notation where the powers are positive. 8 to the power of -6 x 8 to the power of 2 6 to the power of 7 divided by 6 to the power of 12

zepdrix (zepdrix):

So the first one is this, ya? \(\large\rm 5^5\cdot5^5\)

OpenStudy (raphaelll):

Yup thats the first one

zepdrix (zepdrix):

Here is our rule of exponents that will be very helpful to use: \(\large\rm x^a\cdot x^b=x^{a+b}\)

zepdrix (zepdrix):

Notice that when we apply this rule, the `base` stays the same.

zepdrix (zepdrix):

Any ideas how we can use this? :)

OpenStudy (raphaelll):

umm, i just take 5 and add the two powers together?

zepdrix (zepdrix):

\[\large\rm 5^\color{orangered}{5}\cdot5^\color{orangered}{5}=5^{\color{orangered}{5+5}}\]Mmm good!

OpenStudy (raphaelll):

thanks! how do i go about doing the second question since there is a negative power?

zepdrix (zepdrix):

8 to the power of -6 x 8 to the power of 2 \[\large\rm 8^{-6}\cdot8^{2}=8^{-6+2}\]Same thing, addition :) Remember how to add a negative and positive together?

OpenStudy (raphaelll):

oh i see, is it also just addition if division is the middle step?

zepdrix (zepdrix):

division is a little different: \(\large\rm x^a\div x^b=x^{a-b}\)

zepdrix (zepdrix):

6 to the power of 7 divided by 6 to the power of 12 \[\large\rm 6^7\div6^{12}=6^{7-12}\] Make sure you write the 7 and 12 in the correct locations. I was forced to write the 7 first because that shows up on the left of the division. The 7 is smaller, so yes, you should end up with a negative power after you complete the subtraction.

OpenStudy (raphaelll):

okay thank you! so for multiplication it is just adding the two powers, and for division it is subtracting the two powers

zepdrix (zepdrix):

Yes! You might be able to see a connection between the multiplication and division... if you really think about it. Not a big deal though :)

OpenStudy (raphaelll):

you wouldn't happen to know anything about scientific notation would you? it's the last two questions on my assignment that i don't understand

zepdrix (zepdrix):

Can you write out the question? :o That might make it easier for us.

OpenStudy (raphaelll):

26.9 x 463 000 and 55.18 divided by 620

zepdrix (zepdrix):

For the first one, when you do the multiplication you get: \(\large\rm 12,454,700\) This is not scientific notation. For scientific notation, we want to write this as a number... times some power of 10.

zepdrix (zepdrix):

So we want to write the number like this. The number that goes here \(\large\rm \square\) should be a single digit. something larger than 1, and less than 10. So when we look at our number, we can see that THIS is the number that goes in the square: